the perimeter of a standard basketball court is 288 ft. find the dimensions of the basketball court if its length is 44ft. longer than its width.
Found 2 solutions by checkley77, rfer:
Answer by checkley77(12569)

2X+2(X+44)=288
2X+2X+88=288
4X=288-88
4X=200
X=200/4
X=50 FT. IS THE WIDTH.
50+44=94 FT. IS THE LENGTH.
PROOF:
2*50+2*94=288
100+188=288
288=288

Answer by rfer(10432)

width=x=50ft
length=x+44=94ft
-----------
p=2w+2L
2x+2(x+44)=288
2x+2x+88=288
4x=200
x=50